OpenStudy (anonymous):

solve and check 2x^2+7x=9

OpenStudy (anonymous):

This is a quadratic equation, as it has a power of 2 (x^2). You can do this by the following methods: 1) The quadratic formula (long-but-always-work-method) 2) Root method (short-but-might-not-work method) This is what i call it, your book will probably call it something else. Which method do you want me to do for you?

OpenStudy (anonymous):

@Thee121 it doesnt matter i just want 2 learn how to get the right answer

OpenStudy (anonymous):

alright, i will show you the root method then (which is the easiest in my opinion): 1)You put the equation set to 0: \[2x ^{2}+7x-9=0\] Then you identify a, b and c, which are a=2, b=7 and c=-9 (if you want me to explain this to you, tell me) 2) Then you ask yourself: "what two numbers multiplied by each other give you the C, and added together give you B". THIS WILL COME WITH PRACTICE 3)For this question, this cannot be done, as there are no numbers that multiplied together will give you -9 and added together will give you 7. Therefore, you move into the quadratic formula. 4)The quadratic formula is attached to this question. 5)Substitute a, b and c into the equation (which you obtained from step 1) 6) voala, you have the roots! If you have any problems with the quadratic formula, tell me :)

OpenStudy (anonymous):

im having problems @Thee121

OpenStudy (anonymous):

@agent0smith @amistre64 @hartnn @pitamar @mathstudent55

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

Do you still need help?

OpenStudy (texaschic101):

2x^2 + 7x - 9 = 0 a = 2, b = 7, c = -9 x = (-b (+-) sq rt b^2 - 4ac)/2a x = (-7 +- sq rt 7^2 - 4(2)(-9) / 2(2) x = (-7 +- sq rt 49 - 8(-9)) / 4 x = (-7 +- sq rt 49 + 72) / 4 x = (-7 +- sq rt 121)/4 x = (-7 + 11)/4 = 4/4 = 1 x = (-7 - 11)4 = -18/4 = - 9/2 or - 4.5 x = 1 and -4.5 you better get a second opinion on this....I am not real good with the quadratic formula

OpenStudy (anonymous):

@brooklynransome yes i do

OpenStudy (anonymous):

Set the equation equal to zero. \[2x ^{2}+7x−9=0\] Use the quadratic Equation because it always works. \[\frac{ -b \pm \sqrt{b ^{2}- 4ac} }{ 2a }\] Fill in for the variables. \[\frac{ -7 \pm \sqrt{7^{2}- 4 (2)(-9)} }{ 2(2) }\] Simplify: \[\frac{ -7\pm \sqrt{121} }{ 4 }\] \[\frac{ -7 \pm 11 }{ 4 }\]

OpenStudy (anonymous):

Then you do -7 + 11 which = 4. 4/4 = 1. Now you do -7 -11 which = -18. -18 divided by 4 = -4 and 1/9.

OpenStudy (anonymous):

Your two answers are 1 and \[-4\frac{ 1 }{ 9 }\]. Now you have to substitute those answers to check and make sure they are both true.

OpenStudy (texaschic101):

-18 divided by 4 = -4 2/4 = -4 1/2 or - 4.5

OpenStudy (anonymous):

Oh I just did this by hand. THAT MAKES SO MUCH MORE SENSE. haha. so just fill those in and check them.